OpenStudy (anonymous):

Determine the dual basis A={(1,0,0),(1,1,0),(1,1,1)}=(a1,a2,a3) my answer look different to that one in the book but i don't know where did i made the mistake.

OpenStudy (anonymous):

my answer is [1-1 0] ,[1 0 -1 ] and [1 0 0] but they are saying the answer is [1 -1 0],[ 0 1 -1] and [ 0 0 1] help pls!

OpenStudy (anonymous):

@Coolsector help pls

OpenStudy (anonymous):

i am a bit rusty on this, but as much as I remmeber, dual base vectors should satisfy this: \[e_{i}e^{j}=\delta _{i}^{j}\]

OpenStudy (anonymous):

actually i am confused do thet use natural basis or they are using the given vectors to find the answer

OpenStudy (anonymous):

The condition i wrote in my previous message equals to 3 systems of 3 equations in 3 unknowns. Solving each system will give you one vector for the dual base. Here goes the first one which would be for the first dual base vector: 1x+0y+0z= 1 1x+1y+0z=0 1x+1y+1z=0 for second vector: 1x+0y+0z=0 1x+1y+0z=1 1x+1y+1z=0 for third: 1x+0y+0z=0 1x+1y+0z=0 1x+1y+1z=1 solve this 3 systems and you got your vectors

OpenStudy (anonymous):

@REMAINDER